If there is 0.551 mg of As in 348 g of solution, what is the As concentration in ppm?


1
Expert's answer
2021-01-06T09:05:32-0500

No. Of milli moles =0.551/75= 0.0073 mmoles

Molarity = 0.0073/0.348 0.021M